OpenStudy (anonymous):

. Given the linear equation 2y + 6 = 5x, find the y-intercept.

OpenStudy (anonymous):

2y=5x-6 y=2.5x-6 ur y-intercept is -6

OpenStudy (anonymous):

options are 5/2, 3, -3, and -6/5

OpenStudy (callisto):

2y=5x-6 Divide both sides by 2 to get y = mx+c, where c is the y-int. 2y=5x-6 y = 2.5x - 3 What is the y-intercept?

OpenStudy (anonymous):

oh i forgot to divide 6 by 2.... its gonna b -3

OpenStudy (anonymous):

-3?

OpenStudy (anonymous):

srry mady. my bad :|

OpenStudy (callisto):

yup!~